OpenStudy (anonymous):

integral of (3x+1) / ((x^3)(3x+2)(x-1)) dx

OpenStudy (anonymous):

\[\int\limits \frac{ 3x+1 }{ x^3 (3x+2)(x-1)}\]

OpenStudy (anonymous):

I think you use partial fractions here so. \[3x+1=A(x^5)(3x+2)(x-1)+B(x^4)(3x+2)(x-1)\] \[+C(x^3)(3x+2)(x-1)+D(x^6)(x-1)+E(x^6)(3x+2)\]

OpenStudy (anonymous):

using x=1 I was able to get E=4/5 what other numbers for x should I use, did I type this correctly?

OpenStudy (anonymous):

\[\frac{ 3x+1 }{ x^3(3x+2)(x-1) }=\frac{ A }{ x }+\frac{ B }{ x^{2} }+\frac{ C }{ x^{3} }+\frac{ D }{ 3x+2 }+\frac{ E }{ x-1 }\]

OpenStudy (anonymous):

\[3x+1 = Ax^2(3x+2)(x-1) + Bx(3x+2)(x-1) + C(3x+2)(x-1) + Dx^3(x-1) + Ex^3(3x+2)\]

OpenStudy (anonymous):

for A wouldn't you have to use the x^3 though?

OpenStudy (anonymous):

oh, nevermind. I understand XD

OpenStudy (anonymous):

The cut off part: \[Dx^3(x-1) + Ex^3(3x+2)\] No, you do not. You have 5 linear factors on bottom, x*x*x*(3x+2)(x-1). For each of those linear factors, you only need a constant in the each of their fractions for the partial fraction decomp.

OpenStudy (anonymous):

at least that explains the weird answers part XD

OpenStudy (anonymous):

Yeah, you would only need an x^3 on top if you had a factor that was some sort of quartic factor like x^4 + 3x^2 + 1 or something silly like that, lol.

OpenStudy (anonymous):

what numbers for x would you suggest

OpenStudy (anonymous):

0 gave me C=1/2

OpenStudy (anonymous):

what does 1 give then?

OpenStudy (anonymous):

E=4/5

OpenStudy (anonymous):

and then -2/3 for x

OpenStudy (anonymous):

And C should be -1/2, not just 1/2 :)

OpenStudy (anonymous):

D=-81/20

OpenStudy (anonymous):

Sould be -81/40

OpenStudy (anonymous):

haha I meant to type that XD

OpenStudy (anonymous):

So what would we use for A&B since we used up our roots

OpenStudy (anonymous):

We can pick anything we want. Something that will make calculations as easy as possible. In the end, we will need to do a system of equations for A and B. Let's try -1 first and see what we get left with.

OpenStudy (anonymous):

25/8=A+B.... That doesn't look right...

OpenStudy (anonymous):

It still might be. We need a second equation with A and B so we can do a system of equations. So we can pick another x value, like x = 2, and see what we get. Then take those 2 equations to solve for A and B. If the algebra goes wrong and we have issues wih that, there's another way to solve for A and B without picking x-values.

OpenStudy (anonymous):

how would we do that?

OpenStudy (anonymous):

I just need to solve A and B then I know how to do the other stuff